OpenStudy (ohohaye):

Write the equation of the conic section with the given properties A hyperbola with vertices (0,-6) and (0,6) and asymptotes y=3/4x and y=-3/4x

OpenStudy (ohohaye):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

let me think

jimthompson5910 (jim_thompson5910):

how far did you get with this?

OpenStudy (ohohaye):

Nowhere

jimthompson5910 (jim_thompson5910):

what is the midpoint of (0,6) and (0,-6) ?

OpenStudy (ohohaye):

0

OpenStudy (ohohaye):

Or (0,0)

jimthompson5910 (jim_thompson5910):

(0,0) yes

jimthompson5910 (jim_thompson5910):

what is the distance from (0,0) to (0,6)

OpenStudy (ohohaye):

6

jimthompson5910 (jim_thompson5910):

so a = 6 'a' is the distance from the center (0,0) to either vertex

jimthompson5910 (jim_thompson5910):

to find b, we need to solve this equation a/b = 3/4

jimthompson5910 (jim_thompson5910):

a = 6, so a/b = 3/4 6/b = 3/4 what is the value of b?

OpenStudy (ohohaye):

7?

jimthompson5910 (jim_thompson5910):

6/b = 3/4 6*4 = b*3 ... cross multiply 24 = 3b 3b = 24 b = ?????

OpenStudy (ohohaye):

8

jimthompson5910 (jim_thompson5910):

b = 8 is correct

jimthompson5910 (jim_thompson5910):

(h,k) is the center, so (h,k) = (0,0) h = 0 and k = 0

jimthompson5910 (jim_thompson5910):

a = 6 b = 8 h = 0 k = 0 plug those values into \[\Large \frac{(y-k)^2}{a^2} - \frac{(x-h)^2}{b^2} = 1\]

OpenStudy (ohohaye):

\[\frac{ (y-0)^2 }{ 6^2 }-\frac{ (x-0)^2 }{ 8^2 }\]

OpenStudy (ohohaye):

=1

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (ohohaye):

Ok, and would I do the same thing for the Ellipse except with the Ellipse formula?

jimthompson5910 (jim_thompson5910):

ellipses don't have asymptotes though

OpenStudy (ohohaye):

then how would I solve it if it were an Ellipse and it asked for a minor or major axis?

jimthompson5910 (jim_thompson5910):

I'd have to see the full problem

OpenStudy (ohohaye):

Like this: Write the equation of the conic section with the given properties an Ellipse with vertices (0,-5) and (0,5) and a minor axis of length 8

jimthompson5910 (jim_thompson5910):

what is the distance from (0,-5) to (0,5) ?

OpenStudy (ohohaye):

10?

jimthompson5910 (jim_thompson5910):

yes, half of this distance is 5 half of the minor axis (8) is 4 so for \[\Large \frac{(x-h)^2}{a^2} + \frac{(y-k)^2}{b^2} = 1\] we know that h = 0 k = 0 a = 4 b = 5

OpenStudy (ohohaye):

\[\frac{ (x-0)^2 }{ 4^2 }+\frac{ (y-0)^2 }{ 5^2 }=1\]

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (ohohaye):

Ok, thank you

jimthompson5910 (jim_thompson5910):

no problem